OpenStudy (anonymous):

is the answer to: 5x+18y=0 y=5x+18?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

No, it's not. First move y to one side and x to the other: 18y=-5x then you need to divide.

OpenStudy (anonymous):

sorry, typed the question wrong, I ment: 9x+18y = 0

OpenStudy (anonymous):

Okay, so same principle. Just replace the -5 with a -9.

OpenStudy (anonymous):

ok how would you write that? 18y=-9x /9 /9 2y= -1x?

OpenStudy (anonymous):

That is true, but that is not slope intercept form. You would then need to divide both sides by 2, such that y=-1/2x.